Alejandraajtum12:

A 62.0 kg water skier at rest jumps from the dock into a 775 kg boat at rest on the east side of the dock. If the velocity of the skier is 4.50 m/s as she leaves the dock, what is the final velocity of the skier and boat?

peregrine:

m1v1 + m2v2 = (m1 + m2) vf 62 * 4.5 m/s + 775 * 0 = ( 62 + 775) * vf vf = 0.33 m/s east